# LLM-Card: Towards a Description and Landscape of Large Language Models

### v2

@@ 3.1.2 Dependency Tree Construction
3.2 Pipeline
To extract the relationship between LLM names and application scenarios or licences from sentences, a set of rules was developed. These rules identify and extract the relationships between model names, application scenarios, and licences based on the syntactic structure of the sentence. Dependency analysis can not only reveal subject-object relationships but also capture modification relationships and compound structures. As shown in Figure 6, the sentence: GPT-4 enhances Text Generation with more accurate and context-aware outputs, where enhances as a core verb forms the main semantic relationship with the subject GPT-4 and the object Text Generation, while enhances forms the main semantic relationship with GPT-4 and the object Text Generation. Through these parses, we are able to systematically extract and structure this information, providing a basis for more advanced tasks such as information extraction and relational modelling.
Figure 6: Dependency tree for the active voice processing sentence: GPT-4 enhances text generation with more accurate and context-aware outputs.
4 Experiment
@@
4.1 Settings
Complex data formats require complex logic to parse and understand, the Portable Document Format (PDF) is one of the most demanding formats because it is both a data exchange format and a presentation format and has a particularly rigorous tradition of supporting interoperability and consistent presentation Anantharaman et al. (2023). This work mainly uses the PyMuPDF python library to extract the text of PDF. 106 academic papers were processed by defining three dictionaries- LLMs name, licence, and application. 11,051 sentences were extracted through dictionary lookup, and the dataset was constructed through manual review of the final selection of 129 sentences with a link between the name and the licence, and 106 sentences with a link between the model name and the application.
F1 Score is a measure of predictive performance used in this work to evaluate the results of the relation extraction and it is the reconciled mean of Precision and Recall, defined by the formula:
$F1\ Score=2\times\frac{Precision\times Recall}{Precision+Recall}$ (1)
Among them: Precision indicates how many of the samples predicted to be in the positive category are truly positive. The formula is:
$Precision=\frac{TP}{TP+FP}$ (2)
TP: True Positives, samples that the model correctly predicts as positive. FP: False Positives, samples that the model incorrectly predicts as positive. Recall indicates how many of the samples that are true positives are correctly predicted as positives. The formula is:
$Recall=\frac{TP}{TP+FN}$ (3)
@@ 4.1 Settings
Table 1: Example of Relational extraction datasets for large language model and application
@@
Table 2: Example of Relational extraction dataset for large language model and licence
@@ 4.2 Rule Formulation
4.2.1 Active voice processing
@@
Figure 7: Dependency tree for the active voice processing sentence: RoBERTa improves text classification by fine-tuning on large datasets.
Figure 8: Dependency tree for the active voice processing sentence: ERNIE is effective in Named Entity Recognition (NER), particularly in Chinese text.
4.2.2 Passive voice processing
@@
Figure 9: Dependency tree for the passive voice processing sentence XLNet is utilised in Text Generation, providing context-aware sentence completions.
Figure 10: Dependency tree for the passive voice processing sentence: BERT was released under the Apache licence 2.0.
@@ 4.2.3 Evaluation
Table 3: Evaluation metrics for the test dataset
4.3 Visualisation
Figure 11: A visual summary of various large language models, applications, and licences. Blue represents the name of the model, green represents the relationship between the model name and the licence and application entities, and red represents the application and licence of the large language model.
@@ 4.3 Visualisation
5 Conclusions and Future Work
5.1 Conclusions
As a pilot study, the main goal of this work is to validate the basic feasibility of the proposed method LLM-Card, i.e., automatically extract LLMs in published literature and their model cards of model details and their affiliated information. We have successfully constructed two core datasets using dependency parsing and relation extraction techniques: one on the application scenarios of big language models, and the other on the types of licences used in these models. These datasets provide a solid foundation for further research.
In order to further validate the utility of the method, we also conducted real-world testing of the extracted relations and entities using locally deployed large language models. These big language models performed well in handling natural language tasks, especially in the absence of specialised training data, and were able to accurately identify and predict many complex linguistic relations by virtue of their powerful zero-shot learning capabilities. However, despite the model’s good performance in initial tests, there is still room for improvement in its predictions. In order to improve the model’s performance in various complex situations, further instruction training is essential (e.g. methods from Cui et al. (2023); Ren et al. (2024); Micheletti et al. (2024); Belkadi et al. (2024)). Especially when it comes to literature in specialised areas, the model needs targeted training to enhance its understanding of specific terms and complex grammatical structures, thus ensuring the accuracy and applicability of the prediction results.
5.2 Future work
@@
As proof of concept investigation, the scope of validation in this study is limited, focussing on a small number of literature and specific tasks, and has not yet been fully tested on larger datasets. Although the preliminary results show the effectiveness of the method LLM-Card, future research needs to expand the size of the validation dataset and apply more stringent evaluation criteria to comprehensively examine the robustness and applicability of the method.
Ethical Statement
We applied the designed algorithms and methods to publically available literature / scientific articles. It does not release/involve any personal data.
Author Contributions
ST: experimental work and draft first manuscript, LH: co-supervision and revising the manuscript, EMG: co-supervision, GN: PI and co-supervision.
Acknowledgements
LH and GN are grateful to grant support EP/V047949/1 “Integrating hospital outpatient letters into the healthcare data space” (funder: UKRI/EPSRC).

### v3

@@ 3.1.2. Dependency Tree Construction using spaCy
The overall diagram on how we applied dependence parsing rules is presented in Figure 7. A step-by-step interpretation is as follows.
First, we set four initial parameters, subject, object(obj), relation, and is_passive. Through these parameters, we respectively stored the sentence subject information into the subject, the direct object and object part into the obj, the root part between subject and obj as relation, and identified whether the sentence is in passive voice. Secondly, if we identify the sentence as passive voice, we process the passive construction in terms of finding the passive subject, main vern with its auxillaries part, and prepositional object to form the relational triple. If we identify the sentence as the active construction we find the subject part first and check the active complement type in terms of direct object, adjectival, and prepositional respectively to extract direct object(dobj), adjective complement(acomp) with prep phrase, prep phrase with prep object(pobj) to form the triple. Finally, we confirm that the triples are valid. The algorithm illustration using pseudo-code is in Figure 8.
To give interpretation examples, in the left hand of the diagram in Figure 7, for “Check complement type”:
  • 1. Direct object =¿ e.g. “ChatGPT improves NER” =¿ dobj(improves, NER)
  • 2. Adjectival: e.g. “ChatGPT is good at NER” =¿ acomp(is, good); prep phrase: at NER
  • 3. Prepositional: e.g. “ChatGPT applies at NER task” =¿ prep+pobj: at + pobj(NER task)
For more examples, “dobj” is the relation (enhances, Text Generation) in Figure 6; “acomp” is the relation (is, effective), “prep” is the relation (effective, in), and “pobj” is the relation (in, Named Entity Recognition) in Figure 10.
3.2. Rules Applied for Specific Relations
As a pilot study, we first applied rules on LLMs and their (applications, licences). To extract the relationship between LLM names and application scenarios or licences from sentences, a set of rules was developed. These rules identify and extract the relationships between model names, application scenarios, and licences based on the syntactic structure of the sentence. Dependency analysis can not only reveal subject-object relationships but also capture modification relationships and compound structures. As shown in Figure 6, the sentence: GPT-4 enhances Text Generation with more accurate and context-aware outputs, where enhances as a core verb forms the main semantic relationship with the subject GPT-4 and the object Text Generation, while enhances forms the main semantic relationship with GPT-4 and the object Text Generation. Through these parses, we are able to systematically extract and structure this information, providing a basis for more advanced tasks such as information extraction and relational modelling. In the next section, we introduce our experimental settings in detail and the outcomes.
4. Experiment
@@
4.1. Settings
Complex data formats require complex logic to parse and understand, the Portable Document Format (PDF) is one of the most demanding formats because it is both a data exchange format and a presentation format and has a particularly rigorous tradition of supporting interoperability and consistent presentation (Anantharaman et al., 2023). This work mainly uses the PyMuPDF python library to extract the text of PDF. 106 academic papers were processed by defining three dictionaries- LLMs name, licence, and application. 11,051 sentences were extracted through dictionary lookup, and the dataset was constructed through manual review of the final selection of 129 sentences with a link between the name and the licence, and 106 sentences with a link between the model name and the application.
@@ 4.1. Settings
Table 1. Example of Relational extraction datasets for large language model and application
@@
Table 2. Example of Relational extraction dataset for large language model and licence
@@ 4.2. Rule Formulation
4.2.1. Active voice processing
@@
Figure 9. Dependency tree for the active voice processing sentence: RoBERTa improves text classification by fine-tuning on large datasets.
Figure 10. Dependency tree for the active voice processing sentence: ERNIE is effective in Named Entity Recognition (NER), particularly in Chinese text.
4.2.2. Passive voice processing
@@
Figure 11. Dependency tree for the passive voice processing sentence XLNet is utilised in Text Generation, providing context-aware sentence completions.
Figure 12. Dependency tree for the passive voice processing sentence: BERT was released under the Apache licence 2.0.
@@ 4.2.3. Evaluation
Table 3. Evaluation metrics for the test dataset
4.3. Visualisation
Figure 13. A visual summary of various large language models, applications, and licences. Blue represents the name of the model, green represents the relationship between the model name and the licence and application entities, and red represents the application and licence of the large language model.
@@ 4.3. Visualisation
5. Conclusions and Future Work
5.1. Conclusions
As a pilot study, the main goal of this work is to validate the basic feasibility of the proposed method AutoLLM-Card, i.e., automatically extract LLMs in published literature and their model cards of model details and their affiliated information. We have successfully constructed two core datasets using dependency parsing and relation extraction techniques: one on the application scenarios of big language models, and the other on the types of licences used in these models. These datasets provide a solid foundation for further research.
In order to further validate the utility of the method, we also conducted real-world testing of the extracted relations and entities using locally deployed large language models. These big language models performed well in handling natural language tasks, especially in the absence of specialised training data, and were able to accurately identify and predict many complex linguistic relations by virtue of their powerful zero-shot learning capabilities. However, despite the model’s good performance in initial tests, there is still room for improvement in its predictions. In order to improve the model’s performance in various complex situations, further instruction training is essential (e.g. methods from (Cui et al., 2023; Ren et al., 2024; Micheletti et al., 2024; Belkadi et al., 2024)). Especially when it comes to literature in specialised areas, the model needs targeted training to enhance its understanding of specific terms and complex grammatical structures, thus ensuring the accuracy and applicability of the prediction results.
5.2. Future work
@@
We plan to carry more model comparisons with the very recent works such as (Singh et al., 2023; Liu et al., 2024; Yang et al., 2024). We are also implementing new methods regarding our NER component of the AutoLLM-Card pipeline.
@@
As proof of concept investigation, the scope of validation in this study is limited, focusing on a small number of literature and specific tasks, and has not yet been fully tested on larger datasets. Although the preliminary results show the effectiveness of the method AutoLLM-Card, future research needs to expand the size of the validation dataset and apply more stringent evaluation criteria to comprehensively examine the robustness and applicability of the method.
Acknowledgement
This updated version would not have been possible without the scientific and critical review and feedback from Prof Suzan Verbene.
Appendix A Appendix
Figure 14. data cleaning example after filtering out the non-relation entries